OpenStudy (anonymous):

Find all solutions of the given equation csc^2 θ − 4 = 0

OpenStudy (anonymous):

csc^2=4 csc=2

OpenStudy (anonymous):

csc=1/sin=2/1 sin=1/2=pi/6 pi/6+2kpi, 5pi/6+2kpi

OpenStudy (anonymous):

is this correct?